solve the equation by taking square root:
8x^2-3=-155

Answer by RAY100(1637)

8x^2 -3 = -155
.
8x^2 = -152
.
x^2 = -152/8 = - 19
.
taking sqrt of both sides
.
x= sqrt(-19) = 4.36 i,,,,,,,,where i = sqrt(-1)